I have counted the number of statistics problems I do every day. The following data lists...

  1. I have counted the number of statistics problems I do every day. The following data lists the day of the week and how many problems I did on that day in the second column. The third column gives the percent of problems theory suggests I should have done.

Sunday

25

0

Monday

102

20

Tuesday

87

10

Wednesday

122

20

Thursday

76

10

Friday

111

20

Saturday

50

20

Test if I followed the theory at α=10%
Answer #1

Dear your satisfaction is more important, I provided answers with detailed explanation. If you are satisfied with my explanations please THUMBS UP.. Thanks inadvance for your support dear....

This is a chisquare goodness of fit test Ho: Hii Good fit Not a Goodl fit. Test statistic (83) - 6; – E1) Ei follows a chisquTotal questions you have done = 25 +102 +87+122 + 76+llit 50 = 573 Day Expected frequency 0% of 573 = 0 *573 = 0 100 100 SundWe cant have an expected cell with less than 5 et count. one add observations of Sunday and Monday together.Day observed coi) Expected CEi so Cop-Es)?! Sunday + Monday 127 114.6 1 342 Tuesday 187. 57083 15.394 Wednesday 122 114.6 0.